• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 23:54:22 on October 28,1998

    Entry number 3989

    keyword=shift summary

    16:00-18:20 2.030 Gev // (29.3 Mevents)
    18:20-18:40 NMR & field rotation to perendicular
    18:40-21:32 2.030 Gev -| (16.7 Mevents)
    21:32-22:05 NMR &field rotation to perpendicular, Spectrometer change began
    22:05-23:02 EP measurement.
    23:02-mid began 1.873 Gev //

    The 2.030 Gev kinematic was completed. Yves performed an EP measurement
    and M. Finn made several HAPPEX runs per B. Michaels request.
    The beam was very stable throughout the shift with only 4 or 5 very
    short trips. Target NMR gave roughly 35% polarization.